If c 0 +c 1 +c 2 +.........+c n =256,then 2n c 2 is equal to A... 56 B...120 C...28 D...91

If c0+c1+c2+.........+cn=256,then 2ncis equal to
A... 56
B...120
C...28
D...91

Grade:12th pass

1 Answers

Aditya Gupta
5 years ago
we know that c0+c1+c2+.........+cn= 2^n
so, 2^n=256=2^8
or n=8
so, 2nc2 = 16C2= 120, so option B is correct